OpenStudy (anonymous):

Factor completely: x2 − 4x + 9

OpenStudy (diyadiya):

\( Quadratic Formula \) \[x= \frac {-b \pm \sqrt{b^2-4ac}}{2a}\] \( a=1\ b=-4 \ c=9 \) \[x=\frac{4 \pm \sqrt{16-4(1)(9)}}{2} = \frac{4 \pm i2\sqrt{5}}{2}\] \( x=2+i \sqrt{5} \ or \ x=2-i \sqrt{5}\)